OpenStudy (anonymous):

using the substitution method how do i solve y=-x-1 5x-4y=-32

OpenStudy (anonymous):

5x-4(-x-1)=32 Do you need more help? (this is just the substitution)

OpenStudy (anonymous):

would you mind helping me step for step on how to solve it?

OpenStudy (anonymous):

5x−4−x−1=−32 Simplify both sides of the equation. 5x−4−x−1=−32 Simplify 5x−4−x−1=−32 5x+−4+−x+−1=−32 (5x+−x)+(−4+−1)=−32(Combine Like Terms) 4x+−5=−32 4x−5=−32 Add 5 to both sides. 4x−5+5=−32+5 4x=−27 Divide both sides by 4. 4x/4=−27/4 x=−6.75 Hope this helps a bit!